The question states:
A square was made into a rectangle by adding 3 inches to one side, and taking off 3 form the adjacent side. The area of the rectangle is 40 sq.ft. Create an equation in terms of x, to help you solve and find the original area of the square.

Answer by Timnewman(323)

Hi,
Note; square has all sides equal.
let original lenght of the square be "x"
since one side is increased by 3,x will increase by 3
therefore,(x+3)
Also the other is decreased by 3,then,(x-3)
since the area of the rectangle formed is 40,
(x+3)(x-3)=40
x²-3x+3x-9=40
x²-9=40
x²=40+9
x²=49
(This is an equation interms of x)
Hence,
x=\/49
x=7
the lenght of the original square is 7
The area(A) of the original square =7*7
=49ft²
The area of the original square is 49ft²
CHEERS
Timnewman,...